# Downstream Task Guided Masking Learning in Masked Autoencoders Using Multi-Level Optimization

### v1

@@ 4.3 Main Results
Table 2: Test accuracy (%) in linear probing experiments. Results marked with * are taken from AutoMAE (Chen et al., 2023).
@@
ImageNet-1K 55.4 58.5 $*{}^{*}$65.0 $*{}^{*}$68.8 70.2
Linear Probing Results.
Table 2 shows the linear probing results. Our method MLO-MAE demonstrates superior performance compared to baselines across various datasets. Specifically, on ImageNet, MLO-MAE achieves an impressive accuracy of 70.2%, substantially surpassing MAE’s accuracy of 55.4% and U-MAE’s 58.5%. Similarly, on the CIFAR-100 dataset, MLO-MAE continues to outperform, attaining an accuracy of 63.8%, significantly higher than the 46.6% accuracy of MAE and 50.4% accuracy of U-MAE.
Table 3: Accuracy (%) on fine-grained image classification datasets. All methods use ViT-B as the backbone with a patch size of 16. The results of baselines are taken from AutoMAE (Chen et al., 2023). U-MAE did not report results on these three datasets.
@@ 4.4.1 Fine-grained Image Classification
In MLO-MAE, the masking network is trained using a specific downstream classification dataset, raising concerns about potential overfitting and limited generalizability to other datasets. To address this, we performed transfer learning experiments. Given the ViT-B models pretrained on ImageNet using MLO-MAE and baselines and subsequently fine-tuned on ImageNet, we transferred them to perform fine-grained classification by further fine-tuning them on iNaturalist 2019, CUB-200-2011, and Stanford Cars. Classification accuracy results, as presented in Table 3, show that MLO-MAE surpasses all baselines across these datasets. This indicates that the masking network, as learned by MLO-MAE for a particular downstream classification dataset, is capable of generalizing its effectiveness to additional datasets, rather than being overly tailored to that specific downstream classification dataset.
4.4.2 Semantic Segmentation
Table 4: Semantic segmentation results on ADE20K. The results of baselines are taken from AutoMAE (Chen et al., 2023).
Method mIoU
Supervised Pretraining 45.3
MAE 46.1
SemMAE 46.3
AutoMAE 46.4
MLO-MAE (Ours) 49.8
@@ 4.4.2 Semantic Segmentation
We also explored the transferability of the masking network, initially learned through a downstream classification task, to a different task domain: semantic segmentation. Given a ViT-B model pretrained on ImageNet using MLO-MAE or a baseline and subsequently fine-tuned on ImageNet, to transfer it for semantic segmentation, we integrated it as a backbone model into the UPerNet (Xiao et al., 2018) semantic segmentation framework. It was then further fine-tuned on the challenging ADE20K dataset (Zhou et al., 2017) containing 25K images spanning 150 semantic categories. The fine-tuning was conducted by the AdamW optimizer for over 160,000 iterations, with a batch size of 8 and a learning rate of 0.0001.
In Table 4, MLO-MAE showcases a significant improvement over the baselines. Specifically, MLO-MAE attains enhancements in mean Intersection over Union (mIoU) by margins of 3.7% and 3.4% when compared to these baselines. This performance highlights the capability of MLO-MAE in executing dense prediction tasks.
@@ 4.5 Ablation Studies
Figure 2: Comparison of linear probing performances of the BLO-MAE as per Eq.(5) and MLO-MAE on CIFAR-10.
Table 5: Linear probing accuracy (%) on CIFAR-10, under different unrolling steps.
@@
Unrolling Steps.
In this study, we explored how the number of unrolling steps in the Optimization Algorithm, as detailed in Section 3.2, on the final performance. The study was performed on CIFAR-10, with linear probing as the evaluation protocol. Table 5 shows linear probing accuracy under different unrolling steps. The results reveal that an increase in the unrolling steps leads to a gradual improvement in accuracy. This enhancement can be attributed to the fact that a higher number of unrolling steps allows for more frequent updates to the image encoder weights (with more iterations in Stage I), prior to any updates being made to the classification head and masking network. Consequently, this yields a more refined gradient estimation for the parameters of the classification head and masking network, as these estimations are based on the image encoder that has undergone more extensive training. However, it is important to note that increasing the number of unrolling steps also brings in a notable computational overhead. In this context, our observations indicate that while increasing the unrolling step from one to two leads to a 0.5% boost in CIFAR-10 linear probing performance, the gain diminishes to just 0.3% when the unrolling steps are further raised from two to five. This suggests a diminishing return in performance improvement relative to the increased computational demand.
Table 6: Linear probing accuracy (%) on CIFAR-10, under different patch sizes.
@@
We studied how the masking ratio during pretraining affects downstream task performance, using the CIFAR-10 dataset and a linear probing protocol. The results, illustrated in Figure 3, reveal that intermediate masking ratios deliver optimal performance. When the masking ratio is too low, the reconstruction task becomes overly simple, failing to push the image encoder to develop robust representations. Conversely, an excessively high masking ratio makes the task overly challenging, which also impedes the learning of effective representations. Notably, our method demonstrates resilience to changes in masking ratios ranging from 60% to 80%. Within this interval, there is minimal fluctuation in the results of linear probing.
@@
In this study, we investigated the impact of different image patch sizes on the performance of our method when applied to the CIFAR-10 dataset, utilizing the linear probing evaluation protocol. Our experiments focused on three patch sizes: $2\times 2$, $4\times 4$, and $8\times 8$. The results, presented in Table 6, show that the $2\times 2$ patch size achieves a linear probing accuracy of 84.3%, which is 2.4% (absolute) higher than that obtained with the larger $8\times 8$ patch size. These findings suggest that MLO-MAE is more effective when employing a larger number of smaller patches, particularly for small images, such as those in the CIFAR-10 dataset with a size of $32\times 32$. The reason is that smaller patches allow for more precise and detailed candidate masks. This leads to better reconstruction during training and therefore better feature representation learning. Moreover, smaller patches provide a finer grid over the image, allowing the model to capture more detailed and subtle features. This is particularly beneficial for small images, where each pixel can carry significant information.
4.6 Visualizations
To better understand the behavior of the masking network in our MLO-MAE framework, we present a visual illustration in Figure 4, depicting how images are masked at various ratios: 10%, 50%, 75%, and 90%. At a lower masking ratio, such as 10%, our network is more effective in masking the foreground (e.g., the chair), compared to a random masking strategy. As we increase the masking ratio, the attention map reveals that a consistent group of patches in the foreground, which display high attention values (indicated by the yellow area), are selectively masked by our masking network. This is in contrast to random masking, where no specific pattern is observed across different masking ratios. This behavior suggests that our masking network is adept at identifying and masking the informative patches, as opposed to the less informative ones, across a range of masking ratios.
@@ 4.7 Computational Costs
Figure 4: Examples of random masking and MLO-MAE masking applied to the ImageNet-1K validation set at various masking ratios.
@@
ImageNet-1K 1066 hrs 1154 hrs 1344 hrs 1083 hrs
Although MLO-MAE introduces additional computational overhead due to its multi-level optimization, this is balanced by its lower epoch requirement for achieving convergence. In contrast to the 800 epochs needed for standard MAE and SemMAE, MLO-MAE converges in just 50 epochs. This significant reduction in the number of epochs effectively offsets the increased computational demands. The total GPU hours (using Nvidia A100) for MLO-MAE on ImageNet, as shown in Table 7, amount to 1083. This number is comparable to the 1066 hours for MAE and is slightly less than the 1154 hours for SemMAE and the 1344 hours for AutoMAE.
@@ 5 Conclusion
In this paper, we proposed MLO-MAE, a method that automatically learns an optimal masking strategy in Masked Autoencoder (MAE) by leveraging feedback from downstream tasks. Unlike the vanilla MAE which applies uniform patch masking irrespective of their informativeness, MLO-MAE adaptively concentrates on more informative image regions. Different from other MAE methods that do not utilize feedback from downstream tasks for masking-strategy optimization, MLO-MAE uniquely capitalizes on such feedback to refine its masking approach. Our experiments across various datasets demonstrate that MLO-MAE outperforms MAE baselines by learning downstream task guided masking strategies. Furthermore, the representations generated by MLO-MAE exhibit high transferability to a range of downstream tasks, including fine-grained classification and semantic segmentation, highlighting our method’s versatility and effectiveness.
Potential Broader Impact and Ethical Considerations
@@ A.3 Integration of Methods
Implicit differentiation and finite difference approximation are integrated to balance theoretical accuracy with computational feasibility. This combination enhances the robustness and efficiency of our optimization process in the MLO-MAE framework. In the application of these optimization methods, several key considerations are taken into account. Firstly, the choice of $\Delta T$ in the finite difference approximation is a critical factor, as it directly influences the accuracy and stability of the gradient estimation. An appropriate value for $\Delta T$ ensures a balance between precision and numerical stability. Secondly, we address the computational complexity problem inherent in implicit differentiation. This aspect is particularly relevant in deep network architectures, where computational resources can be a limiting factor. To mitigate this, we optimize the use of implicit differentiation to balance computational demands with the need for accurate gradient computation. Lastly, maintaining numerical stability is paramount in both methods. Techniques such as gradient normalization and careful arithmetic handling are employed to ensure that the computations remain stable, especially in scenarios where small numerical errors can significantly impact the overall results.
This comprehensive optimization algorithm is pivotal in enabling the efficient training and validation of the MLO-MAE framework.
@@ A.4 Differentiability of MLO-MAE Framework
To further address the non-differentiability issue in our MLO-MAE framework incurred by mask selection, the SoftSort (Prillo & Eisenschlos, 2020) technique—a differentiable approximation of sorting operations—replacing the conventional, non-differentiable ”argsort” operation, can be integrated. SoftSort enables the learning of a continuous relaxation of sorting, vital for gradient-based optimization and for crafting nuanced, performance-enhancing masks. This represents a significant leap in refining the effectiveness of the reconstruction process. However, given the computational demands of SoftSort, we propose a simplified approach that approximates the base Jacobian with an identity matrix similar to  (Choe et al., 2023), thereby simplifying the gradient during backpropagation and enhancing the efficiency by ”jumping over” the non-differentiable argsort operation. By treating non-differentiable operations as identity functions during backpropagation, this method allows gradients from the reconstruction loss to flow as if the masking operations were inherently differentiable. This strategy significantly speeds up training by enabling the use of standard gradient-based optimization techniques. In our MLO-MAE, the masking network employs soft masks, which, through a sigmoid activation function $\sigma(.)$, assign each image patch a continuous value from 0 to 1. This assignment reflects the probability of masking, informed by both masked and unmasked patches. By approximating the gradients for non-differentiable functions as identity functions during backpropagation, our method enables the differentiation of the MAE reconstruction loss, $\mathcal{L}_{rec}$, relative to the masking network parameters $T$. This differentiation is facilitated by incorporating the output probabilities of $\sigma(.)$ into the reconstruction loss $\mathcal{L}_{rec}$, as detailed in Equation (1).

### v2

@@ 4.2 Main results
Table 2: Test accuracy (%) in linear probing experiments.
@@
ImageNet-1K 68.0 58.5 65.0 68.8 70.2
Table 3: Accuracy (%) on fine-grained image classification datasets. All methods use ViT-B as the backbone with a patch size of 16.
@@
Table 4: Semantic segmentation results on ADE20K.
Method mIoU
Supervised Pretraining 45.3
MAE 48.1
SemMAE 46.3
AutoMAE 46.4
MLO-MAE (Ours) 49.8
Linear probing results.
Table 2 shows the linear probing results. Our method MLO-MAE demonstrates superior performance compared to baselines across various datasets. Specifically, on ImageNet, MLO-MAE achieves an accuracy of 70.2%, substantially surpassing MAE’s accuracy of 55.4% and U-MAE’s 58.5%. Similarly, on the CIFAR-100 dataset, MLO-MAE continues to outperform, attaining an accuracy of 63.8%, significantly higher than the 46.6% accuracy of MAE and 50.4% accuracy of U-MAE.
@@ 4.3.1 Fine-grained image classification
In MLO-MAE, the masking network is trained using a specific downstream classification dataset, raising concerns about potential overfitting and limited generalizability to other datasets. To address this, we performed transfer learning experiments on labeled fine-grained classification datasets including iNaturalist 2019, CUB-200-2011, and Stanford Cars. Classification accuracy results, as presented in Table 4, show that MLO-MAE surpasses all baselines across these datasets. This indicates that the masking network, as learned by MLO-MAE for a particular downstream classification dataset, is capable of generalizing its effectiveness to additional datasets, rather than being overly tailored to that specific downstream classification dataset.
4.3.2 Semantic segmentation and object detection
@@
We also explored the transferability of the masking network, initially learned through a downstream classification task, to other tasks including semantic segmentation and object detection. Given a ViT-B model pretrained on ImageNet using MLO-MAE or a baseline and subsequently fine-tuned on ImageNet, to transfer it for semantic segmentation, we integrated it as a backbone model into the UPerNet (Xiao et al., 2018) semantic segmentation framework. It was then further fine-tuned on the challenging ADE20K dataset (Zhou et al., 2017) containing 25K images spanning 150 semantic categories. The fine-tuning was conducted by the AdamW optimizer for over 160,000 iterations, with a batch size of 8 and a learning rate of 0.0001. In Table 4, MLO-MAE showcases a significant improvement over the baselines. Specifically, MLO-MAE attains enhancements in mean Intersection over Union (mIoU) by margins of 3.7% and 3.4% when compared to these baselines. This performance highlights the capability of MLO-MAE in executing dense prediction tasks.
Table 5: Object detection result of MLO-MAE and baselines on MS-COCO detection task.
Method AutoMAE MAE MLO-MAE
$\text{AP}^{\text{box}}$ (%) 50.5 50.3 51.1
Following MAE, we also adapt MLO-MAE pretrained ViT-B model for the use of an FPN backbone in Mask R-CNN. As shown in Table 5, MLO-MAE pretrained backbone model performs better than all baselines (51.1 comparing to 50.5 and 50.3, $\text{AP}^{\text{box}}$). We did not include SemMAE and U-MAE as they did not report on MS-COCO detection. Due to space limits, we defer the results of object detection on PASCAL VOC 2007 to Appendix D.1.
4.4 Continued pretraining
Table 6: Continued pretraining on PDDB and PAD-UFES. All settings are initialized with weights of ViT-B pretrtained by MAE on ImageNet. Runtime measured in GPU hours on A100.
PDDB PAD-UFES
Method Acc(%) Runtime Acc(%) Runtime
No continued pretraining 88.6 2132 75.0 2132
MAE continued pretraining 89.3 2132 + 39 75.4 2132 + 3
MLO-MAE pretraining (no IN pretrain) 90.2 2132 + 37 75.9 2132 + 3
MLO-MAE continued pretraining 92.7 2132 + 37 77.6 2132 + 3
We further investigated the effectiveness of MLO-MAE in a continued pretraining setting. Starting with a ViT-B model pretrained by MAE on ImageNet, we applied MLO-MAE pretraining to 2 datasets, PDDB (Barbedo et al., 2018) and PAD-UFES (Pacheco et al., 2020). Table 6 compares the test accuracy (%) across three settings: (1) no continued pretraining, where the model is directly fine-tuned using labels; (2) continued pretraining on target dataset using MAE, followed by fine-tuning; (3) MLO-MAE pretraining solely on target dataset, followed by fine-tuning; and (4) continued pretraining on target dataset using MLO-MAE, followed by fine-tuning. Our results show that continued pretraining with MLO-MAE significantly outperforms both MAE-based pretraining and no pretraining in both datasets. This highlights the practical advantage of MLO-MAE: by leveraging MAE pretraining once, subsequent tasks can benefit from fast, efficient continued pretraining with MLO-MAE, delivering substantial performance gains without requiring the computationally expensive process of large-scale pretraining from scratch for each downstream task.
@@ 4.5 Ablation studies
Unrolling steps.
In this study, we explored how the number of unrolling steps in the Optimization Algorithm, as detailed in Section 3.2, affects the final performance. The study was performed on CIFAR-10, with linear probing as the evaluation protocol. Table 8 shows linear probing accuracy under different unrolling steps. The results reveal that an increase in the unrolling steps leads to a gradual improvement in accuracy. This enhancement can be attributed to the fact that a higher number of unrolling steps allows for more frequent updates to the image encoder weights (with more iterations in Stage I), prior to any updates being made to the classification head and masking network. Consequently, this yields a more refined gradient estimation for the parameters of the classification head and masking network, as these estimations are based on the image encoder that has undergone more extensive training. However, it is important to note that increasing the number of unrolling steps also brings in a notable computational overhead. In this context, our observations indicate that while increasing the unrolling step from one to two leads to a 0.5% boost in CIFAR-10 linear probing performance, the gain diminishes to just 0.3% when the unrolling steps are further raised from two to five. This suggests a diminishing return in performance improvement relative to the increased computational demand.
Table 7: Linear probing accuracy (%) on CIFAR-10 under different unrolling steps.
@@
Table 8: Linear probing accuracy (%) on CIFAR-10 under different patch sizes.
@@
We studied how the masking ratio during pretraining affects downstream task performance, using the CIFAR-10 dataset and a linear probing protocol. The results, illustrated in Figure 3, reveal that intermediate masking ratios deliver optimal performance. When the masking ratio is too low, the reconstruction task becomes overly simple, failing to push the image encoder to develop robust representations. Conversely, an excessively high masking ratio makes the task overly challenging, which also impedes the learning of effective representations. Notably, our method demonstrates resilience to changes in masking ratios ranging from 60% to 80%. Within this interval, there is minimal fluctuation in the results of linear probing. Additional experiment on dynamic mask ratio can be found in Appendix D.4.
@@
In this study, we investigated the impact of different image patch sizes on the performance of our method when applied to the CIFAR-10 dataset, utilizing the linear probing evaluation protocol. Our experiments focused on three patch sizes: $2\times 2$, $4\times 4$, and $8\times 8$. The results, presented in Table 8, show that the $2\times 2$ patch size achieves a linear probing accuracy of 84.3%, which is 2.4% (absolute) higher than that obtained with the larger $8\times 8$ patch size. These findings suggest that MLO-MAE is more effective when employing a larger number of smaller patches, particularly for small images, such as those in the CIFAR-10 dataset with a size of $32\times 32$. The reason is that smaller patches allow for more precise and detailed candidate masks. This leads to better feature representation learning. Moreover, smaller patches provide a finer grid over the image, allowing the model to capture more detailed and subtle features. This is particularly beneficial for small images, where each pixel can carry significant information.
Semantic segmentation as downstream feedback.
Table 9: Mean intersection over Union (MIoU) in using semantic segmentation as the downstream task on ADE20K
Method (mIoU) MAE MLO-MAE (Ours)
ADE20K 28.8 29.1
In the base MLO-MAE framework, image classification was used as the downstream task to guide the masking network. In this section, we evaluate the generalizability of MLO-MAE by adapting semantic segmentation as the downstream task in Stages II and III. Specifically, we adopted the ADE20K dataset for both pretraining and semantic segmentation evaluation. The results, summarized in Table 9, demonstrate that when semantic segmentation is used as the downstream feedback, MLO-MAE consistently outperforms the baseline. This highlights that the effectiveness of MLO-MAE is not limited to a specific downstream task, underscoring its versatile and robust design.
@@ 4.6 Computational costs
ImageNet-1K 2132 hrs 1154 hrs 1344 hrs 1083 hrs
Although MLO-MAE introduces additional computational overhead due to its multi-level optimization, this is balanced by its lower epoch requirement for achieving convergence. In contrast to the 800 epochs needed for standard MAE and SemMAE, MLO-MAE converges in just 50 epochs. This significant reduction in the number of epochs effectively offsets the increased computational demands. The total GPU hours (on Nvidia A100 GPU) for MLO-MAE on ImageNet, as shown in Table 10, amount to 1083. This number is less than those of baselines while our method achieves better test accuracy than baselines as shown in Tables 1 and 2.
@@ 5 Conclusion
In this paper, we proposed MLO-MAE, a method that automatically learns an optimal masking strategy in Masked Autoencoder (MAE) by leveraging feedback from downstream tasks. Unlike the vanilla MAE which applies uniform patch masking irrespective of their informativeness, MLO-MAE adaptively concentrates on more informative image regions. Different from other MAE methods that do not utilize feedback from downstream tasks for masking-strategy optimization, MLO-MAE uniquely capitalizes on such feedback to refine its masking approach. Our experiments across various datasets demonstrate that MLO-MAE outperforms MAE baselines by learning downstream task guided masking strategies. Furthermore, the representations generated by MLO-MAE exhibit high transferability to a range of downstream tasks, including fine-grained classification, semantic segmentation, and object detection, highlighting our method’s versatility and effectiveness.
6 Broader impact
@@
7 Limitations
Our multi-level optimization framework introduces two main limitations. First, the hypergradient computation incurs significant computational overhead. We mitigate this by using gradient approximation, where we control the approximation level through the number of unrolling steps. This trade-off balances computational cost with optimization accuracy. Second, our approach increases GPU memory usage due to the inclusion of a classification head and the need to load three separate training datasets simultaneously. Although our current implementation does not explicitly address this, techniques like gradient accumulation could reduce memory demands. Future work can focus on developing more efficient multi-level optimization algorithms to systematically overcome these limitations.
Acknowledgments
This research was supported by NSF IIS2405974 and NSF IIS2339216.
@@ A.3 Integration of Methods
Implicit differentiation and finite difference approximation are integrated to balance theoretical accuracy with computational feasibility. This combination enhances the robustness and efficiency of our optimization process in the MLO-MAE framework. In the application of these optimization methods, several key considerations are taken into account. Firstly, the choice of $\Delta T$ in the finite difference approximation is a critical factor, as it directly influences the accuracy and stability of the gradient estimation. An appropriate value for $\Delta T$ ensures a balance between precision and numerical stability. Secondly, we address the computational complexity problem inherent in implicit differentiation. This aspect is particularly relevant in deep network architectures, where computational resources can be a limiting factor. To mitigate this, we optimize the use of implicit differentiation to balance computational demands with the need for accurate gradient computation. Lastly, maintaining numerical stability is paramount in both methods. Techniques such as gradient normalization and careful arithmetic handling are employed to ensure that the computations remain stable, especially in scenarios where small numerical errors can significantly impact the overall results. This comprehensive optimization algorithm is pivotal in enabling the efficient training and validation of the MLO-MAE framework.
Algorithm 1 MLO-MAE Optimization Algorithm
1:Input: Training dataset $\mathcal{D}_{tr}$, validation dataset $\mathcal{D}_{val}$
2:Output: Optimized parameters $E^{*}(T)$, $D^{*}(T)$, $C^{*}(E^{*}(T))$, $T^{*}$
3:procedure MAE Pretraining($\mathcal{D}_{u},T$)
4:     Initialize encoder $E$ and decoder $D$
5:     for each image $X_{i}\in\mathcal{D}_{u}$ do
6:         Patchify image into $\bar{X}_{i}$
7:         Compute masked image using $\sigma(\mathcal{M}_{j}(X;T,r),X;T)$
8:         Update $E$, $D$ by minimizing $\sum\limits_{X\in\mathcal{D}_{u}}\sum\limits_{j=1}^{N\times r}\sigma(\mathcal{M}_{j}(X;T,r),X;T)\mathcal{L}_{rec}(X-\mathcal{M}(X;T,r),\mathcal{M}_{j}(X;T,r);E,D)$
9:     end for
10:     return $E^{*}(T)$, $D^{*}(T)$
11:end procedure
12:procedure Classification Head Training($E^{*}(T),\mathcal{D}_{tr}$)
13:     Freeze $E^{*}(T)$, initialize classifier $C$
14:     for each image $X_{i}\in\mathcal{D}_{tr}$ do
15:         Update $C$ by minimizing $\mathcal{L}_{cls}(\mathcal{D}_{tr};E^{*}(T),C)$
16:     end for
17:     return $C^{*}(E^{*}(T))$
18:end procedure
19:procedure Validation Optimization($E^{*}(T),C^{*}(E^{*}(T)),\mathcal{D}_{val}$)
20:     Freeze $E^{*}(T)$, $C^{*}(E^{*}(T))$
21:     Optimize $T$ by minimizing $\mathcal{L}_{cls}(\mathcal{D}_{val};E^{*}(T),C^{*}(E^{*}(T)))$
22:     return $T^{*}$
23:end procedure
@@ A.4 Differentiability of MLO-MAE Framework
To further address the non-differentiability issue in our MLO-MAE framework incurred by mask selection, the SoftSort (Prillo & Eisenschlos, 2020) technique—a differentiable approximation of sorting operations—replacing the conventional, non-differentiable "argsort" operation, can be integrated. SoftSort enables the learning of a continuous relaxation of sorting, vital for gradient-based optimization and for crafting nuanced, performance-enhancing masks. This represents a significant leap in refining the effectiveness of the reconstruction process. However, given the computational demands of SoftSort, we propose a simplified approach that approximates the base Jacobian with an identity matrix similar to  (Choe et al., 2023), thereby simplifying the gradient during backpropagation and enhancing the efficiency by "jumping over" the non-differentiable argsort operation. By treating non-differentiable operations as identity functions during backpropagation, this method allows gradients from the reconstruction loss to flow as if the masking operations were inherently differentiable. This strategy significantly speeds up training by enabling the use of standard gradient-based optimization techniques. In our MLO-MAE, the masking network employs soft masks, which, through a sigmoid activation function $\sigma(.)$, assign each image patch a continuous value from 0 to 1. This assignment reflects the probability of masking, informed by both masked and unmasked patches. By approximating the gradients for non-differentiable functions as identity functions during backpropagation, our method enables the differentiation of the MAE reconstruction loss, $\mathcal{L}_{rec}$, relative to the masking network parameters $T$. This differentiation is facilitated by incorporating the output probabilities of $\sigma(.)$ into the reconstruction loss $\mathcal{L}_{rec}$, as detailed in Equation (1).
A.5 Convergence Property of the MLO-MAE framework
Previous research has extensively analyzed the convergence properties of optimization algorithm for bi-level optimization (BLO) problems. For instance,  Ji et al. (2021) establish sharper non-asymptotic convergence rates for implicit differentiation (AID) and iterative differentiation (ITD) methods, addressing both deterministic and stochastic bilevel problems with improved computational complexity.  Grazzi et al. (2020) analyze the iteration complexity of hypergradient computation, providing explicit non-asymptotic convergence bounds and demonstrating the efficiency of AID with conjugate gradient under contraction assumptions. Bao et al. (2021) focus on the stability and convergence of bilevel optimization by deriving expectation bounds for the generalization gap, emphasizing how regularization strategies improve convergence. Bao et al. (2021) address bilevel problems with nonconvex lower-level structures, introducing the IAPTT-GM framework, which ensures convergence even in the absence of lower-level convexity. Zhang et al. (2021) propose a convergence-guaranteed approach for bilevel optimization using stochastic implicit gradients, ensuring stationary point convergence while reducing computational overhead.
While the analyses in the aforementioned works are primarily on bi-level optimization problems, the convergence proof can be readily extended to the multi-level cases. In addition our emipirical results demonstrate that MLO-MAE framework consistently converges and outperforms the baseline methods as shown in Section 4. The training loss curves in Fig 4 also corroborate with our empirical convergence finding.
(a) Masking Loss
(b) Finetune Loss
(c) Pretrain Loss
Figure 4: Loss curves for masking, fine-tuning, and pretraining.